OpenStudy (anonymous):

A ball is released from a hot air balloon moving downward with a velocity of -10.0 meters/second and a height of 1,000 meters. How long did it take the ball to reach the surface of Earth? Given: g = -9.8 meters/second2.

OpenStudy (irishboy123):

using \(x = ut + \dfrac{1}{2}a t^2 \) will give you a quadratic to solve. u = 10, a = 9.8 alternatively, calculate velocity v when it hits ground from \(v^2 = u^2 + 2ax\) and use \(v = u + at\) to get t.